# Generalized Bayesian deep reinforcement learning

### v1

@@ Abstract
Bayesian reinforcement learning (BRL) is a method that merges principles from Bayesian statistics and reinforcement learning to make optimal decisions in uncertain environments. Similar to other model-based RL approaches, it involves two key components: (1) Inferring the posterior distribution of the data generating process (DGP) modeling the true environment and (2) policy learning using the learned posterior. We propose to model the dynamics of the unknown environment through deep generative models assuming Markov dependence. In absence of likelihood functions for these models we train them by learning a generalized predictive-sequential (or prequential) scoring rule (SR) posterior. We use sequential Monte Carlo (SMC) samplers to draw samples from this generalized Bayesian posterior distribution. In conjunction, to achieve scalability in the high dimensional parameter space of the neural networks, we use the gradient based Markov chain Monte Carlo (MCMC) kernels within SMC. To justify the use of the prequential scoring rule posterior we prove a Bernstein-von Misses type theorem. For policy learning, we propose expected Thompson sampling (ETS) to learn the optimal policy by maximising the expected value function with respect to the posterior distribution. This improves upon traditional Thompson sampling (TS) and its extensions which utilize only one sample drawn from the posterior distribution. This improvement is studied both theoretically and using simulation studies assuming discrete action and state-space. Finally we successfully extend our setup for a challenging problem with continuous action space without theoretical guarantees.
@@ 1 Introduction
Effective learning and decision-making within perpetually changing dynamic systems are imperative for applications like controlling automated machinery and enabling robotic navigation. Reinforcement learning (RL) stands out as a potent tool in these domains, enabling agents to acquire knowledge through trial and error and responses from the environment. Its versatility has resulted in its widespread utilization across a range of sectors, encompassing automated vehicles (Guan et al., 2020), robotics (Kormushev et al., 2013), healthcare (Yu et al., 2021), finance (Deng et al., 2016), and various natural language processing (NLP) applications (Uc-Cetina et al., 2023), recommender systems (Chen et al., 2023) and so on.
@@
Typically, an RL task can be expressed using a Markov decision process (MDP) (Sutton and Barto, 2018) framework. MDPs are a classical framework for sequential decision-making, where a policy maps the state space to the action space. The multitude of RL algorithms existing in the literature can be broadly segregated into two categories: those reliant on a model (‘model-based’ algorithms) and those that facilitate policy learning in a ‘model-free’ manner. Model-free algorithms like Q-learning algorithms (Clifton and Laber, 2020) and policy gradient methods (Sutton and Barto, 2018) directly learn from the history generated by real time interactions. Conversely, model-based approaches like dynamic programming (Sutton and Barto, 2018), Monte Carlo tree search (Coulom, 2006), PILCO (Deisenroth and Rasmussen, 2011) etc. use knowledge about the environment to design a policy.
When an efficient model of the environment is available or can be learned, model-based RL (Moerland et al., 2023) strategies are more sample-efficient than model-free methods. In practical scenarios, explicit models for environment dynamics are often unavailable, but simulator models may be accessible. In the recent advances of model-based RL, neural network based predictive models (Nagabandi et al., 2018), (Kaiser et al., 2019); conditional GANs (Charlesworth and Montana, 2020), (Zhao et al., 2021) have been successfully adopted to model the complex dynamical systems underlying the environment. These models are extremely potent as seen in a multitude of recent applications, but their complexity makes training challenging due to the lack of a tractable objective function. Hence previous work either approximates a divergence function (as in adversarial training regime (Goodfellow et al., 2020) or the Fisher score using a function class parametrized by neural networks (Gurney, 2018). Both of these strategies are sample-greedy and computationally expensive defeating the purpose of online RL, further the adversarial training is known to be notoriously difficult to train and faces issues (e.g. mode collapse (Arora et al., 2018)) due to the accumulation of the bias in each iteration of training.
For more robust training of the underlying model, probabilistic model predictive control (MPC) (Kamthe and Deisenroth, 2018), conditional variational inference (VI) to capture multimodal stochasticity (Moerland et al., 2017), and Bayesian neural networks (BNNs) (Depeweg et al., 2016) etc have been explored in the literature. Notably, the Bayesian approach offers the additional benefit of quantifying uncertainty in the model’s learning of the environment’s dynamics. However, in real-world scenarios, the likelihood function required for these updates is often intractable, necessitating likelihood-free inference.
In this work we mainly focus on Bayesian model-based RL approach which integrates Bayesian inference with RL to make decisions in uncertain environments. It comprises two primary components: (1) Bayesian inference of the underlying MDP that governs environment dynamics, and (2) designing a policy based on the learned model. Hence, our main contribution is in providing a likelihood-free framework for training Bayesian generative NN to model the underlying MDP and a strategy to use this model for policy learning.
Bayesian inference of the MDP
To perform likelihood-free inference, we define a generalized Bayesian scoring rule posterior (Pacchiardi et al., 2024b) using predictive-sequential (or prequential) scoring rules (Dawid, 1984; Dawid and Vovk, 1999; Dawid, 1991). Prequential SRs evaluate the fit of observed data to predictions conditioned on the past data using strictly proper scoring rules (e.g. energy score, kernel score etc., see (Gneiting and Raftery, 2007) for a full review). Prequential SR has been previously used to train generative network based models for probabilistic forecasting (Pacchiardi et al., 2024a). The same idea can be extended to define a generalized posterior which uses the prequential SR as an alternative for the negative log likelihood of a model. We further show that a Bernstein-von Mises type theorem (showing contraction of the posterior to the true parameter value) holds for these generalized posteriors when a deep generative MDP is used to model the environment. Even in the absence of a tractable likelihood function, an unbiased estimate of the SR can be derived using the observed data and samples from the generative models. Further, when the generative model is differentiable, an unbiased estimator of the gradient of the generalized posterior is easy to derive which enables the exploration of the parameter space in an efficient and scalable manner.
Markov Chain Monte Carlo (MCMC) methods are classical tools for sampling from a posterior distribution. Since the observed data in reinforcement learning (RL) scenarios is typically acquired sequentially through interactions with the environment, we opt for SMC samplers (Del Moral et al., 2006) to infer the posterior distribution of model parameters. To enhance the performance of SMC samplers, particularly in the high-dimensional parameter spaces of the generative networks, we incorporate gradient-based MCMC kernels to move the particles. By computing an unbiased estimator of the gradient of the prequential scoring rule, we can leverage the extensive literature on stochastic gradient MCMC methods (Nemeth and Fearnhead, 2021). Specifically, we utilize preconditioned adapted stochastic gradient Langevin dynamics (SGLD) (Jones and Leimkuhler, 2011; Patterson and Teh, 2013; Chen et al., 2016), which are well-suited for exploring high-dimensional parameter spaces. However, when the simulator model is non-differentiable, a common challenge in many RL problems, gradient-based kernels become problematic. In such cases, we can employ zeroth-order gradient estimates (Liu et al., 2020), which rely solely on simulations from the generative models, though this may introduce some estimation bias and variance.
Policy search
Bayesian RL using Thompson sampling (Russo et al., 2018) and their extensions (Osband et al., 2013) are known to have excellent performance with good regret bounds, but while training the policy to explore the parameter space they only utilize one sample drawn from the posterior distribution. However, there is sparse evidence in the literature (Dimitrakakis and Ortner, 2022) showing improved performance whilst using more samples from the inferred posterior distribution. Building upon this, we propose a new policy learning methodology called expected Thompson sampling (ETS) which learns the optimal policy by maximizing the expected value function w.r.t. the inferred posterior distribution. In practice, we estimate the expected value function using samples drawn from the posterior. We also show both using simulation based and theoretical studies that the regret of the policy learned via ETS decreases as the number of posterior samples increases.
@@
Therefore, our proposed approach combines the robustness of scoring rules for Bayesian inference, employs gradient-based SMC for high dimensional parameter spaces, and provides a strategy to design a policy, by potentially better handling model uncertainty. The main contributions of our work are as follows:
  • A likelihood-free framework is proposed to enable the ETS for MDPs without likelihoods (eg. deep generative models) and we show that it is more sample-efficient compared to model-free methods.
  • We provide a theoretical bound for the approximation error in the action value function due to using ETS for generalized posterior.
  • While ETS-based policy learning is computationally intensive due to posterior sampling of model parameters, we make it feasible and scalable by utilizing the SMC for the sequential learning task in hand and gradient based MCMC kernels to handle high-dimensional parameter space.
@@
Figure 1: Bayesian deep RL: The above diagram shows an episodic posterior update followed by a policy update routine for $k$th episode ($k=1,2,\ldots$). Let us assume the episode length to be $\tau$, and $\pi_{0}({\theta})$ to be the prior on the model parameters. Prequential Scoring Rule ($\mathcal{S}$) is used to compute the generalized posterior which is based on true interaction data(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X}$) and a simulation(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X(\theta)}$) from the deep generative model $\mathcal{M}(\theta)$. We obtain the samples, $\{\theta_{1},\theta_{2},\ldots,\theta_{n}\}$ from the generalized posterior ($\pi_{k}(\theta)\propto\pi_{k-1}(\theta)\exp(-\mathcal{S}(X^{(k)},X(\theta)))$) via Sequential Monte Carlo (SMC) samplers. These samples are used to simulate $n$ trajectories of interaction (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X(\theta_{1}),X(\theta_{2}),\ldots,X(\theta_{n})}$) from the model $\mathcal{M}(\theta)$. Optimal policy is trained by maximizing the averaged action value function computed from the $n$ simulated trajectories. The new policy, $\mu^{(k+1)}$ is then used to interact with the true Environment in the next episode.
@@
In Section 2 we provide an overview of Bayesian model-based RL and the relevance of Thompson sampling. In Section 2.1 we introduce expected Thompson sampling (ETS) and demonstrate its application on a simple “chain-task” example for which the likelihood function is available. Section 3 details the generalized Bayesian inference framework, with Section 3.2 discussing the properties of the prequential scoring rule posterior and Section 3.3 explaining the SMC sampler used to sample from the posterior. In Section 4, we derive an error bound for the approximate action-value function under the ETS framework. Simulation studies are presented in Section 5 to evaluate the approach, and Section 6 summarizes key findings and conclusions.
2 Bayesian Model-based Reinforcement learning
@@
In RL, an agent (eg. a video game player) interacts with an unknown environment (eg. the virtual reality inside the game) by taking actions causing the transition of the environment to a new state, yielding rewards in the process. Suppose, at time $t$, the environment was observed at the state $s_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the state space. After the agent’s action $a_{t}\in\mathcal{A}$ (the action space $\mathcal{A}$), the environment moves to the next state, $s_{t+1}\in\mathcal{S}$, and the agent receives a reward $r_{t+1}\in\mathcal{R}$ where $\mathcal{R}$ is the reward space. The goal of the agent is to devise a strategy that selects actions based on the current state to maximize cumulative rewards. Under the assumption that the new state ($s_{t+1}$) only depends upon the previous state and action ($s_{t},a_{t}$) and not on $\{s_{u},a_{u}:u<t\}$ (Markovian assumption), the environment can be considered as a Markov decision process (MDP), as defined below.
@@
A Markov decision process (MDP) $\mathbb{M}$ on the state space $\mathcal{S}$ and action space $\mathcal{A}$, can be defined by the distribution of the initial state $\rho$ and the transition probabilities $P(S_{t+1}=s_{t+1}|S_{t}=s_{t},A_{t}=a_{t})$ [^1] Here, uppercase letters denote random variables, with their lowercase versions representing their realizations. and $P(R_{t+1}=r_{t+1}|S_{t}=s_{t},A_{t}=a_{t})$. The ultimate objective is to develop a policy $\mu$ that maximizes the total reward over the long term in the future, where $\mu:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping of the state space to the the action space. This involves making decisions—choosing actions based on the current state—to navigate the environment in a way that accrues the maximum possible reward. Consequently, it becomes an optimization problem where the objective function is dependent on future rewards. To formalize this, we derive the value of a policy $\mu$ as the expected discounted return from any initial state if the policy $\mu$ is followed thereafter to interact with the environment. Hence, for any time $t$, we define the discounted return as $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}{R}_{t+k+1}$. Then the value function of a policy $\mu$ at the state $S_{t}=s$, is defined as,
@@
Calculation of the value function is not straightforward since it involves taking expectations with respect to the MDP, which is not known in advance. Common approaches include using Bellman equations to approximate the value function based on observations from the environment received through interactions. These approaches are often sample-inefficient, as precise approximation of expectations requires extensive training data. To mitigate this sample inefficiency, in model based RL a parametric model for the MDP is assumed. If we can learn a good model emulating the true MDP, this can solve the above issue as we can draw many samples from the parametric MDP from which the optimal policy can be learned. But this efficiency can only be achieved if we can learn a good model even with smaller training datasets, which we answer positively here. A parametric MDP $\mathbb{M}_{\theta}$, parameterized by $\theta=(\theta_{st},\theta_{r})\in\Theta$, is defined over the state space, $\mathcal{S}$ and action space $\mathcal{A}$. We further assume the action space being finite, meaning $\mathcal{A}$ is countable and $|\mathcal{A}|=n_{a}<\infty$. [^2] Our proposed method and theoretical work will be developed under this assumption, but we will show empirically an extension of this for continuous action space, leaving the theoretical development as future work. $\mathbb{M}_{\theta}$ can be described by the distribution of the initial state $\rho$ and
@@
The distributions $P_{\theta_{st}}$ and $P_{\theta_{r}}$ parameterized by $\theta_{st}$ and $\theta_{r}$ model the state transitions and reward distributions, respectively, for $\mathbb{M}_{\theta}$. For simplicity, the value function and the action-value function of a policy $\mu$ under $\mathbb{M}_{\theta}$ will be denoted as $V^{\theta}_{\mu}$ and $Q^{\theta}_{\mu}$ respectively . Let =$\mathcal{M}=\{\mathbb{M}_{\theta}:\theta\in\Theta\}$ represent a class of Markov decision processes parameterized by $\theta$. If the true underlying MDP, denoted by $\mathbb{M}_{0}$, does or does not belong to the model class $\mathcal{M}$, then we call $\mathbb{M}_{0}$ being well-specified or misspecified respectively.
@@
Assume we have generated a trajectory of state-action-reward-next state sequences of length $T$ through interacting with the unknown environment $\mathbb{M}$, ${H}_{T}=\{(s_{t},a_{t},r_{t+1},s_{t+1})\}_{t=1}^{T}$. To learn $\mathbb{M}_{\theta}$ or the value of the parameters $\theta$ from ${H}_{T}$, we would first define the log likelihood function for $\theta$ with respect to the observed trajectory under $\mathbb{M}_{\theta}$ as,
@@
Following Bayesian inference, we would update prior beliefs about the model parameters with the likelihood function using Bayes theorem
@@
Bayesian model-based RL uses the posterior distribution of parametric MDP $\mathbb{M}_{\theta}$ learnt via Bayesian inference, to learn an optimal policy under the unknown uncertain environments. We first introduce Thompson sampling, a popular approach which only uses one parametric MDP sampled from the posterior distribution, and then propose expected Thompson sampling in Section 2.1 which instead uses many independent and identically drawn parametric MDPs from the posterior distribution.
@@
A Bayesian solution to this challenge begins with a prior belief on the mean rewards, representing our initial knowledge about the environment. As the agent interacts with the environment, it collects a sequence of action-reward pairs. This information is used to update the prior, refining the posterior distribution of the rewards. The posterior distribution is used to design a policy based on this updated knowledge.
In TS, the agent draws a sample from the updated posterior of the reward distribution and chooses the arm associated with the highest sampled reward to interact with the environment in the next round. Thus, TS offers an effective method address the exploration-exploitation dilemma in MAB problems. This approach has been generalized to reinforcement learning (RL) tasks as well (Gopalan and Mannor, 2015; Ouyang et al., 2017). In RL contexts, alongside rewards, the agent observes the environment’s state. In this case, the posterior distribution of the MDP parameters, $\theta$ becomes the quantity of interest.

### v2

@@ Abstract
Bayesian reinforcement learning (BRL) is a method that merges principles from Bayesian statistics and reinforcement learning to make optimal decisions in uncertain environments. As a model-based RL method, it has two key components: (1) inferring the posterior distribution of the model for the data-generating process (DGP) and (2) policy learning using the learned posterior. We propose to model the dynamics of the unknown environment through deep generative models, assuming Markov dependence. In the absence of likelihood functions for these models, we train them by learning a generalized predictive-sequential (or prequential) scoring rule (SR) posterior. We used sequential Monte Carlo (SMC) samplers to draw samples from this generalized Bayesian posterior distribution. In conjunction, to achieve scalability in the high-dimensional parameter space of the neural networks, we use the gradient-based Markov kernels within SMC. To justify the use of the prequential scoring rule posterior, we prove a Bernstein-von Mises-type theorem. For policy learning, we propose expected Thompson sampling (ETS) to learn the optimal policy by maximising the expected value function with respect to the posterior distribution. This improves upon traditional Thompson sampling (TS) and its extensions, which utilize only one sample drawn from the posterior distribution. This improvement is studied both theoretically and using simulation studies, assuming a discrete action space. Finally, we successfully extended our setup for a challenging problem with a continuous action space without theoretical guarantees.
@@ 1 Introduction
Effective learning and decision-making within perpetually changing dynamic systems are crucial for applications like controlling automated machinery and enabling robotic navigation. Reinforcement learning (RL) stands out as a potent tool in these domains, allowing the agents to acquire knowledge through trial and error and responses from the environment. Its versatility has resulted in its widespread utilization in a variety of sectors, including automated vehicles (Guan et al., 2020), robotics (Kormushev et al., 2013), healthcare (Yu et al., 2021), finance (Deng et al., 2016), various applications of natural language processing (NLP) (Uc-Cetina et al., 2023), recommendation systems (Chen et al., 2023), and so on.
@@
Under the classical framework, an RL task can be expressed using a Markov decision process (MDP) (Sutton and Barto, 2018). The multitude of RL algorithms existing in the literature can be broadly segregated into two categories: those reliant on a model (‘model-based’ algorithms) and those that facilitate policy learning in a ‘model-free’ manner. Model-free algorithms like Q-learning algorithms (Clifton and Laber, 2020) and policy gradient methods (Sutton and Barto, 2018) learn directly from the history generated by real-time interactions. In contrast, model-based approaches such as dynamic programming (Sutton and Barto, 2018), Monte Carlo tree search (Coulom, 2006), PILCO (Deisenroth and Rasmussen, 2011) etc. use knowledge about the environment to design a policy.
When a reliable model of the environment is available or can be learned, model-based RL approaches (Moerland et al., 2023) tend to be significantly more sample efficient than model-free methods. In this paradigm, Thompson sampling (TS) (Russo et al., 2018), also known as posterior sampling for RL (PSRL) (Strens, 2000), offers an effective solution to the exploration-exploitation trade-off and enjoys strong theoretical guarantees, including provable regret bounds (Osband et al., 2013; Ouyang et al., 2017). However, its application has been largely restricted to settings with simple MDPs with a tractable likelihood function of the parameters of the model, allowing posterior inference over model parameters.
Recent advances in model-based RL have leveraged powerful neural network-based predictive models (Nagabandi et al., 2018; Kaiser et al., 2019) and conditional GANs (Charlesworth and Montana, 2020; Zhao et al., 2021) to capture complex environment dynamics. Although highly expressive, these models often lack a tractable likelihood function, making training and inference challenging. Previous work addresses this by approximating divergences (e.g., through adversarial training (Goodfellow et al., 2020)) or using surrogate scores such as Fisher information (Gurney, 2018). However, the absence of a well-defined likelihood hinders posterior inference, limiting the applicability of Thompson sampling and its Bayesian extensions.
Our main aim in this work is to provide a tool to perform Thompson sampling for model-based RL when the models considered are deep generative models. Traditional likelihood-free methods such as approximate Bayesian computation (ABC) have been used for TS in this context (Dimitrakakis and Tziortziotis, 2013), but their scalability is limited due to the curse of dimensionality. An alternative is to use scoring rules (SR)(Gneiting and Raftery, 2007), which allow inference when the model lacks a tractable likelihood but is easy to simulate from. Recently, predictive-sequential or prequential scoring rules (Dawid, 1984; Dawid and Vovk, 1999) have shown promise in the training of deep generative models for forecasting tasks (Pacchiardi et al., 2024a). Building on this, we construct a generalized posterior (Bissiri et al., 2016) using the prequential score as a surrogate for the log-likelihood of MDPs modeled by generative networks. TS can then be performed using this generalized posterior. In particular, we employ sequential Monte Carlo (SMC) (Del Moral et al., 2006) samplers enhanced with preconditioned gradient-based Markov kernels Chen et al. (2016) to efficiently explore and sample from the posterior distributions.
Thompson sampling (TS) is a simple yet effective policy learning strategy that relies on a single sample from the posterior over model parameters. Although some recent work (Dimitrakakis and Ortner, 2022) suggests that the use of multiple samples can improve performance, this is still underexplored. Motivated by this, we propose expected Thompson sampling (ETS), a policy learning approach that optimizes the expected action-value function in the posterior distribution. In practice, this expectation is approximated using multiple posterior samples. Through both theoretical analysis and simulations, we demonstrate that the regret of ETS-based policies decreases as the number of samples increases.
@@
Therefore, our proposed approach combines the robustness of scoring rules for Bayesian inference, employs SMC for high dimensional parameter spaces, and provides a strategy to design a policy, by potentially better handling model uncertainty. The main contributions of our work are as follows:
  • A scalable, likelihood-free TS framework for deep generative models in RL.
  • A theoretical bound on the value function approximation error under ETS.
  • An efficient SMC-based implementation of ETS that scales to high-dimensional settings.
@@
Figure 1: Generalized Bayesian deep RL: The diagram illustrates the episodic posterior and policy update process for the $k$th episode ($k=1,2,\ldots$) with episode length $\tau$. Starting with a prior $\pi_{0}(\theta)$ on model parameters, the generalized prequential posterior $\pi_{k}(\theta)\propto\pi_{k-1}(\theta)\exp(-\mathcal{S}(X^{(k)},X(\theta)))$ is computed using a scoring rule $\mathcal{S}$ based on real interaction data $X$ and model simulations $X(\theta)$. SMC is used to draw posterior samples ${\theta_{1},\ldots,\theta_{n}}$, which generate simulated trajectories from $\mathcal{M}(\theta)$. An optimal policy is then trained by maximizing the expected action value over these $n$ trajectories, and the updated policy $\mu^{(k+1)}$ is used in the next episode.
@@
In Section 2 we provide an overview of model-based Bayesian RL and the relevance of Thompson sampling. In Section 2.1 we introduce expected Thompson sampling (ETS) and demonstrate its application on a simple “chain-task” example for which the likelihood function is available. Section 3 details the generalized Bayesian inference framework, with Section 3.2 discussing the properties of the prequential scoring rule posterior and Section 3.3 explaining the SMC sampler used to sample from the posterior. In Section 4, we derive an error bound for the approximate action-value function under the ETS framework. Simulation studies are presented in Section 5 to evaluate the approach, and Section 6 summarizes key findings and conclusions. All proofs of the lemmas and theorems, as well as implementation details of the algorithms, are provided in the supplementary material.
2 Model-based Bayesian Reinforcement learning
@@
In RL, an agent (eg. a video game player) interacts with an unknown environment (eg. the virtual reality inside the game) by taking actions that cause the transition of the environment to a new state, yielding rewards in the process. Suppose, at time $t$, the environment was observed at the state $s_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the state space. After the agent’s action $a_{t}\in\mathcal{A}$ (the action space $\mathcal{A}$), the environment moves to the next state, $s_{t+1}\in\mathcal{S}$, and the agent receives a reward $r_{t+1}\in\mathcal{R}$ where $\mathcal{R}$ is the reward space. The goal of the agent is to devise a strategy that selects actions based on the current state to maximize cumulative rewards. Under the assumption that the new state ($s_{t+1}$) only depends upon the previous state and action ($s_{t},a_{t}$) and not on $\{s_{u},a_{u}:u<t\}$ (Markovian assumption), the environment can be considered as a Markov decision process (MDP), as defined below.
@@
A Markov decision process (MDP) $\mathbb{M}$ on the state space $\mathcal{S}$ and action space $\mathcal{A}$, can be defined by the distribution of the initial state $\rho$ and the transition probabilities $P(S_{t+1}=s_{t+1}|S_{t}=s_{t},A_{t}=a_{t})$ [^1] Here, uppercase letters denote random variables, with their lowercase versions representing their realizations. and $P(R_{t+1}=r_{t+1}|S_{t}=s_{t},A_{t}=a_{t})$. The ultimate objective is to develop a policy $\mu$ that maximizes the total reward over the long term in the future, where $\mu:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping of the state space to the action space. This involves making decisions—choosing actions based on the current state—to navigate the environment in a way that accrues the maximum possible reward. Consequently, it becomes an optimization problem where the objective function is dependent on future rewards. To formalize this, we derive the value of a policy $\mu$ as the expected discounted return from any initial state if the policy $\mu$ is followed thereafter to interact with the environment. Hence, for any time $t$, we define the discounted return as $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}{R}_{t+k+1}$. Then the value function of a policy $\mu$ at the state $S_{t}=s$, is defined as,
@@
Calculation of the value function is not straightforward since it involves taking expectations with respect to the MDP, which is not known in advance. Common approaches include using Bellman equations to approximate the value function based on observations from the environment received through interactions. These approaches are often sample-inefficient, as a precise approximation of expectations requires extensive training data. To mitigate this sample inefficiency, in model based RL a parametric model for the MDP is assumed. If we can learn a good model emulating the true MDP, this can solve the above issue as we can draw many samples from the parametric MDP from which the optimal policy can be learned. But this efficiency can only be achieved if we can learn a good model even with smaller training datasets, which we answer positively here. A parametric MDP $\mathbb{M}_{\theta}$, parameterized by $\theta=(\theta_{st},\theta_{r})\in\Theta$, is defined over the state space, $\mathcal{S}$ and action space $\mathcal{A}$. We further assume the action space be finite, meaning $\mathcal{A}$ is countable and $|\mathcal{A}|=n_{a}<\infty$. [^2] Our proposed method and theoretical work will be developed under this assumption, but we will show empirically an extension of this for continuous action space, leaving the theoretical development as future work. $\mathbb{M}_{\theta}$ can be described by the distribution of the initial state $\rho$ and
@@
The distributions $P_{\theta_{st}}$ and $P_{\theta_{r}}$ parameterized by $\theta_{st}$ and $\theta_{r}$ model the state transitions and reward distributions, respectively, for $\mathbb{M}_{\theta}$. For simplicity, the value function and the action-value function (or, the $Q$-function) of a policy $\mu$ under $\mathbb{M}_{\theta}$ will be denoted as $V^{\theta}_{\mu}$ and $Q^{\theta}_{\mu}$ respectively. Let =$\mathcal{M}=\{\mathbb{M}_{\theta}:\theta\in\Theta\}$ represent a class of Markov decision processes parameterized by $\theta$. If the true underlying MDP, denoted by $\mathbb{M}_{0}$, does or does not belong to the model class $\mathcal{M}$, then we call $\mathbb{M}_{0}$ being well-specified or misspecified respectively.
@@
Suppose we have generated a trajectory of state-action-reward-next state sequences of length $T$ through interacting with the unknown environment $\mathbb{M}$, given by $H_{T}={h}_{T}$ where $h_{T}=\{(s_{t},a_{t},r_{t+1},s_{t+1})\}_{t=1}^{T}$. To learn $\mathbb{M}_{\theta}$ or the value of the parameters $\theta$, we first define the log likelihood function for $\theta$ with respect to the observed trajectory under $\mathbb{M}_{\theta}$ as,
@@
We can update prior beliefs about the model parameters with the likelihood function using the Bayes theorem
@@
model-based Bayesian RL uses the posterior distribution of parametric MDP $\mathbb{M}_{\theta}$ learned via Bayesian inference, to learn an optimal policy under unknown uncertain environments. We first introduce Thompson sampling, a popular approach that only uses one parametric MDP sampled from the posterior distribution, and then propose expected Thompson sampling in Section 2.1 which instead uses many independent and identically drawn parametric MDPs from the posterior distribution.
@@
A Bayesian solution to this challenge begins with a prior belief in the mean rewards, representing our initial knowledge about the environment. As the agent interacts with the environment, it collects a sequence of action-reward pairs. This information is used to update the prior, refining the posterior distribution of the rewards. The posterior distribution is used to design a policy based on this updated knowledge.
In TS, the agent draws a sample from the updated posterior of the reward distribution and chooses the arm associated with the highest sampled reward to interact with the environment in the next round. Thus, TS offers an effective method to address the exploration-exploitation dilemma in MAB problems. This approach has been generalized to reinforcement learning (RL) tasks as well (Gopalan and Mannor, 2015; Ouyang et al., 2017). In RL contexts, alongside rewards, the agent observes the environment’s state. In this case, the posterior distribution of the MDP parameters, $\theta$ becomes the quantity of interest.